OpenStudy (anonymous):

find the vertex for y=2(x-12) squared + 3

OpenStudy (shubhamsrg):

=> (y-3) = 4(1/2)(x-12)^2 now compare this with y = 4ax^2 had it been y=4ax^2 , vertex would have been (0,0) now, had it been (y-3) = 4ax^2 , the whole graph would have been shifted 3 units up, so new vertex would have been (0,3) but it is (y-3) = 4a(x-12)^2 , so whole graph is shifted 12 units the right. so new vertex = (12,3)